# Millen Kanabar and Michael Gastpar School of Computer and Communication Sciences, EPFL

### v3

@@ I-A Contributions
We provide the first (to the best of our knowledge) information-theoretic lower bounds on the $\ell_{2}^{2}$ and $\ell_{1}$ losses at each stage in terms of properties of the estimates at previous stages in Theorem 1 (for the data accumulation workflow) and Corollary 4 (for the data replacement workflow). We also propose a sequence of estimators that is order-optimal under certain regimes with stage-wise performance guarantees in Theorem 2. We show conditions where there are gaps between the upper bound and lower bounds on the minimax loss and the oracle-assisted minimax loss the data accumulation workflow (Propositions 2 and 3) and provide examples where these conditions are satisfied. Finally, we use our lower bound for the data replacement workflow to find a lower bound on the worst-case loss given a time-independent upper bound.
We find that the minimax loss stays within a constant factor of the oracle-assisted one when $\alpha_{t}$ is lower bounded by a constant. When $\alpha_{t}$ decays to 00, i.e. when synthetic sampling outpaces real sampling, the effective sample size of batch $t$ is proportional to $n_{t}\alpha_{t}^{2}$ (as opposed to simply $n_{t}\alpha_{t}$ for the oracle-assisted setting) in most practically conceivable conditions. As a consequence, the ratio of the minimax loss to the oracle-assisted minimax loss can grow unbounded with $t$ (e.g. Claim 1). We specify the conditions for this regime in Proposition 2; the basic requirement is that $\alpha_{t}$ does not decay too quickly.
@@
$\alpha_{i}=\Omega(1)$ (Proposition 1) Both minimax losses are $\Theta\left(\frac{1}{\sum_{i=0}^{t}n_{i}\alpha_{i}}\right)$
$\alpha_{t}\downarrow 0$, but $\alpha_{t}=\Omega\left(\frac{1}{\sum_{i=0}^{t-1}n_{i}\alpha_{i}^{2}}\right)$ (Proposition 2) Bounds match, minimax loss is $\Theta\left(\frac{1}{\sum_{i=0}^{t}n_{i}\alpha_{i}^{2}}\right)$ $\Theta\left(\frac{1}{\sum_{i=0}^{t}n_{i}\alpha_{i}}\right)$
@@
A few additional technical conditions are required for the results shown in the table: The lower bound on the loss at stage $t$ in Proposition 3 assumes the existence of estimators matching the lower bounds up to a universal constant factor for stages 00 through $t-1$ and the lower bounds in general only hold when $\sum_{i=0}^{t}n_{i}\alpha_{i}^{2}+h_{t}(\nicefrac{{1}}{{4k}})\geq k/4$. An elementary upper bound when the hypothesis conditions for Theorem 2 are not met is given in Corollary 3.
@@ I-B Related work
I-C Notation
@@ II Problem Setup
The joint distribution of the random vector $(X_{0}^{n_{0}},\dots,X_{t}^{n_{t}})$ is denoted as $\bar{\mathbf{P}}_{t,({p,\{n_{i},\alpha_{i}:i\geq 0\}})}$. We drop the sample sizes and sampling probabilities from the subscript when they are obvious from the context. By convention, $\alpha_{0}=1$.
The definition of this sequence encapsulates the following process: at stage 00, an estimate $\hat{P}_{0}$ is produced using a batch of $n_{0}$ i.i.d. samples from $p$. At every subsequent stage $t+1$, a batch of $n_{t+1}$ samples is produced, with each sample being distributed as $p$ with probability $\alpha_{t+1}$ and as $\hat{P}_{t}$ with probability $1-\alpha_{t+1}$. Since this process can be seen as conditional sampling based on the outcome of a biased coin-flip, we refer to the outcome of this flip as the ‘identity’ of the source.
@@ III-A Bounds on the minimax loss
The lower bounds on the minimax loss at stage $t$ are given in terms of the performance of the previous estimates, encapsulated in the error term $h_{t}(\epsilon)$, defined as follows:
Definition 3.
The error term $h_{t}(\epsilon)$ is defined as
$h_{t}(\epsilon):=\sum_{i=0}^{t}n_{i}\alpha_{i}g_{i}(\epsilon),$
where
$g_{i}(\epsilon)$ $:=\sup_{p}\max_{j}\bar{\mathbf{P}}_{i-1,({p})}\left\{\hat{P}_{i-1}[j]<p[j]-\epsilon\right\}$ $i\neq 0$
denotes the worst case (over component and distribution) probability of an $\epsilon$-error in $\hat{P}_{i-1}$ under $\bar{\mathbf{P}}_{i-1,({p})}$, and $g_{0}(\epsilon)=0\ \forall\epsilon$.
We are now ready to state the main results of this paper.
@@
whenever $\sum\limits_{i\in[0:t]}n_{i}\alpha_{i}^{2}\,+h_{t}(\epsilon)\geq k/4$.
@@
We can directly observe that when the error term $h_{t}(\nicefrac{{1}}{{4k}})$ is small, the minimax errors are within constant factors of each other; and, as we will see in the baseline, the erasure of the source information leads to a penalty of $\alpha_{t}$ on the effective sample size of batch $t$ (as compared to when the estimators know which distribution each sample is drawn from). As a corollary of the analysis of Theorem 2, we also describe the worst-case error when the sequence of empirical estimators without further processing is used. This can be used in place of Theorem 2 if the hypothesis conditions are not satisfied, and might also be of independent interest.
@@
We note that this bound increases with $t$, and therefore it is perhaps useful only to use the upper bound $\nicefrac{{1}}{{n_{0}}}$ obtained by using the order-optimal estimator for just the $0^{\mathrm{th}}$ batch. However, it is interesting to note that in both cases, there exists a constant upper bound independent of $t$—the sum in the RHS of (5) is exactly $\nicefrac{{1}}{{n_{0}}}$ plus a (lower) Riemann sum of the function $\nicefrac{{1}}{{x^{2}}}$ from $n_{0}$ to $\sum_{i=0}^{t}n_{i}$ with intervals $n_{i}$. Therefore, the upper bound on the minimax error stays $\mathcal{O}(1)$. In fact, if each batch is the same size $n$, we find an upper bound of $\nicefrac{{\pi^{2}}}{{6}}\cdot\nicefrac{{1}}{{n}}$, recovering the additional penalty factor of $\nicefrac{{\pi^{2}}}{{6}}$ highlighted by [12, 16].
@@ III-B Baseline: oracle-assisted minimax loss
To understand the impact of not knowing the source identity on the minimax loss, it is useful to compare the upper and lower bounds presented above to the loss in the case where the identity of the source of each sample is known in the form of an auxiliary random vector $W:=W_{i}^{n_{i}}\sim(\mathrm{Bern}(\alpha_{i}))^{n_{i}},i\in\{1,\dots,t\}$ provided to the estimator by an oracle such that, conditioned on whether $W_{i}[s]$ is 00 or $1$, the sample $X_{i}[s]$ is distributed either as $p$ or $\hat{P}_{i-1}$ respectively.
@@ IV Discussion
It is evident from the bounds shown above that if the mixing factor $\alpha_{t}$ is constant (or even $\Omega(1)$), we can show that the base and oracle-assisted minimax losses are within constant factors of each other; the case where the probability converges to zero is more interesting. This corresponds to cases where the rate of synthetic sampling outpaces the real sampling rate by a considerable margin. Such cases might arise when it is significantly cheaper to sample the synthetic distribution, perhaps also exacerbated by the fact that the estimates get closer to the real distribution with every new batch. The behavior of the bounds in such cases depends on the rates of growth of $\sum_{i=0}^{t}n_{i}\alpha_{i}$ and $\sum_{i=0}^{t}n_{i}\alpha_{i}^{2}$. The absence of source information at the estimator leads to a penalty factor of roughly $\max(\alpha_{i},g_{i-1}(\nicefrac{{1}}{{4k}}))$ in the ‘effective sample size’ of each batch; we show that this can lead to large significant gaps between the base and oracle-assisted losses. Depending on whether or not the $\alpha_{i}$’s are comparable to the $g_{i-1}(\nicefrac{{1}}{{4k}})$’s, there might also be a gap between the guarantees provided by the upper and the lower bounds. In the sequel, we comment on the impact of the error term on the behavior of the bounds and the restrictions on the order-optimal estimator, show cases where differences in the aforementioned rates of growth imply gaps between the base and oracle-assisted losses, and conclude with a summary of the regimes where it is possible to comment precisely on the minimax error.
@@ IV-A The error term $h_{t}({\nicefrac{{1}}{{4k}}})$
The multi-stage lower bounds depend on the performance of estimators in previous stages through the probabilities $g_{i}\left(\nicefrac{{1}}{{4k}}\right)$. We note that the error term is small compared to the other term in the denominator, $\sum_{i=0}^{t}n_{i}\alpha_{i}^{2}$, in quite a lot of cases. Using Chebyshev’s inequality (as demonstrated in the upcoming examples) leads to an upper bound on each of these probabilities in terms of the minimax error $r_{i-1,k}^{\ell}$; the upper and lower bounds thus match whenever $n_{i}\alpha_{i}^{2}$’s are larger than or comparable to $1/\sum_{j=0}^{i-1}n_{j}\alpha_{j}^{2}$. We also note that it is necessary to use Chebyshev’s inequality—as opposed to stronger concentration results such as Hoeffding-style bounds that rely on the specific properties, such as subgaussianity in the case of the Hoeffding bound, of the estimators—to derive a universal upper bound on these probabilities for any sequence of estimators that achieves order-optimal performance.
Intuitively, the quantities $g_{i}(\nicefrac{{1}}{{4k}})$ capture the worst-case deviation probability of any component of $\hat{P}_{i-1}$ from $p$ by more than $\nicefrac{{1}}{{4k}}$. Since we focus consider cases where $\hat{P}_{t}$ converges to $p$, this probability is expected to decay to zero. It is perhaps surprising, then, to also observe that the lower bounds decrease as the probabilities $g_{i}(\nicefrac{{1}}{{4k}})$ increase—suggesting that it might be possible to improve the performance of future estimators by estimating badly in the present. The source of this apparent improvement is the loosening of the upper bound on the KL divergence in (16) in instances when the previous estimators perform badly. This suggests that it might be easier for an estimator to perform well when the synthetic samples are from a distribution sufficiently removed from the real one; the sequence of estimators in Theorem 2 does not share this behavior.
@@ IV-B The order-optimal estimator
The sequence of estimators in Theorem 2 do not always return distributions; indeed, an assumption on the sample sizes and mixing factors is required to make sure that the outputs stay in $\Delta_{k}$ at every stage. Each estimate in the sequence is also kept unbiased and uncorrelated from the previous estimate; this helps keep the analysis from becoming almost prohibitively cumbersome. A careful analysis of a more general class of nonlinear estimators might help bridge the gap with the lower bounds, proving the error term necessary, or drop the assumptions required. Alternatively, a tighter analysis bounding the aforementioned divergence might lead to a lower bound that matches the upper bound wherever valid. For the rest of this paper, we analyze the lower bounds assuming ‘good’ estimators—estimators with worst-case loss within at most a constant factor of the lower bound—exist, and are chosen at every stage. In quite a few cases, as it turns out, we find that the error term is small as compared to the rest of the denominator, and the analysis is tight if optimal or near-optimal estimators are assumed to have been used. We now show the main results in action via two examples, one where the bounds match and the error term can be ignored, and one where the error term dominates, but still leads to a gap w.r.t. the oracle-assisted loss.

### v4

@@ I-A Contributions
We provide the first (to the best of our knowledge) information-theoretic lower bounds on the $\ell_{2}^{2}$ and $\ell_{1}$ losses at each stage in terms of properties of the estimates at previous stages in Theorem 1 (for the data accumulation workflow) and Corollary 4 (for the data replacement workflow). We also propose a sequence of estimators that is order-optimal under certain regimes with stage-wise performance guarantees in Theorem 2. We show conditions where there are gaps between the upper bound and lower bounds on the minimax loss and the oracle-assisted minimax loss the data accumulation workflow (Propositions 2 and 3) and provide examples where these conditions are satisfied. We briefly discuss the performance of estimators that do not account for the introduction of synthetic samples and show that results shown in the Model Collapse literature, where collapse is avoided by simply following the accumulate workflow still hold in our setting (Corollary 3). Finally, we use our lower bound for the data replacement workflow to find a time-independent lower bound on the worst-case (in ground truth distribution) loss given a time-independent upper bound.
We find that the minimax loss stays within a constant factor of the oracle-assisted one when $\alpha_{t}$ is lower bounded by a constant. When $\alpha_{t}$ decays to 0, i.e. when synthetic sampling outpaces real sampling, the effective sample size of batch $t$ is proportional to $n_{t}\alpha_{t}^{2}$ (as opposed to simply $n_{t}\alpha_{t}$ for the oracle-assisted setting) in most practically conceivable conditions. As a consequence, the ratio of the minimax loss to the oracle-assisted minimax loss can grow unbounded with $t$ (e.g. Claim 1). We specify the conditions for this regime in Proposition 2; the basic requirement is that $\alpha_{t}$ does not decay too quickly.
@@
$\alpha_{i}=\Omega(1)$ - (Proposition 1) Both base and oracle-assisted minimax losses are $\Theta\left(\frac{1}{\sum_{i=0}^{t}n_{i}\alpha_{i}}\right)$
$\alpha_{t}\downarrow 0$, but $\alpha_{t}=\Omega\left(\frac{1}{\sum_{i=0}^{t-1}n_{i}\alpha_{i}^{2}}\right)$ - (Proposition 2) Bounds match, minimax loss is $\Theta\left(\frac{1}{\sum_{i=0}^{t}n_{i}\alpha_{i}^{2}}\right)$ $\Theta\left(\frac{1}{\sum_{i=0}^{t}n_{i}\alpha_{i}}\right)$
@@
A few additional technical assumptions are required for the results shown in the table: these are discussed in the corresponding statements.
@@ I-B Related work
A central thread in these works is that the distribution of the data used in later generations of estimates (or training later generations of models) is sampled from functions of the empirical distributions of data seen by earlier generations. This is explicit in the works studying generative machine learning settings, and implicit in works studying model collapse in settings like, for example, regression. While our focus in this work is on the evolution of explicit estimates of the true distribution itself, we believe the results hold relevance for most settings considered in the literature.
We note, however, that our setting differs from these works in our assumption that the estimator can account for the presence of unlabeled synthetic data in the corpus of samples. In contrast, the above works study estimators that act on (treated or untreated) mixed datasets no differently than they would on real datasets. Since our setting also allows for such estimators, lower bounds presented here also function as lower bounds for settings considered in previous model collapse literature.
Our setting also differs from the various settings in the literature where some of the synthetic samples are drawn from estimates, perhaps of a different distribution than the true distribution, based on independent, unseen data 111e.g. in [21], where a careful weighting of these ‘surrogate’ synthetic samples proves beneficial. In our setting, all information about the true distribution available to the estimators is contained in the samples that they observe—minimax rates using labeled synthetic samples from (semi-)independent estimates remain open.
I-C When does a model collapse?
There is no current consensus in on what precisely constitutes model collapse [^2] A list of definitions of model collapse and further discussion can be found in [22].. However, based on the above results, we make the following observations vis-à-vis the various definitions in literature:
I-C1 ‘Non-collapse’
This work mainly focuses on cases where enough real samples are available such that the estimates eventually converge to the true distribution, or remain within a small radius around it. This avoids, by assumption, the conditions for model collapse (collapsing variance, unbounded growth of the expected loss etc.) discussed in prior work. The primary difference is perhaps a focus in the early literature on the replace workflow, where sequences of estimates behave essentially like unreinforced random walks. Even so, our analysis of the replace workflow provides an impossibility result for the expected loss, showing that even the best-case loss cannot be too far from the average-case loss in the oracle-annotated setting.
I-C2 Change in scaling laws with (eventually) ubiquitous synthetic data
When the mixing factor $\alpha_{t}$, i.e. the probability of sampling the true distribution decreases to 0, no matter how slowly, we can show gaps between the minimax loss with and without the presence of synthetic samples. This points to a change in scaling laws, and hence does constitute model collapse according to the criteria in [14]. This is a result of the sampling process itself, and cannot be mitigated without further knowledge of the true distribution, e.g. thorough a (possibly noisy) oracle or via human feedback. In this sense, the introduction of synthetic samples changes the fundamental hardness of the estimation problems regardless of the estimation procedures used.
I-C3 Unchanged scaling in the presence of persistent fractions of real samples
We also find that such gaps do not exist when $\alpha_{t}$ is is bounded below by a positive constant, i.e. when there is a non-vanishing probability of drawing a real sample in each batch. In such cases, model collapse in the discrete distribution estimation setting can be avoided entirely in all accepted senses of the term.
I-D Notation
@@ II Problem Setup
The joint distribution of the random vector $(X_{0}^{n_{0}},\dots,X_{t}^{n_{t}})$ is denoted as $\bar{\mathbf{P}}_{t,({p,\{n_{i},\alpha_{i}:i\geq 0\}})}$. We drop the sample sizes $n_{t}$ and mixing probabilities $\alpha_{t}$ from the subscript when they are obvious from the context. By definition, $\alpha_{0}=1$.
The definition of this sequence encapsulates the following process: at stage 0, an estimate $\hat{P}_{0}$ is produced using a batch of $n_{0}$ i.i.d. samples from $p$. At every subsequent stage $t+1$, a batch of $n_{t+1}$ samples is produced, with each sample being distributed as $p$ with probability $\alpha_{t+1}$ and as $\hat{P}_{t}$ with probability $1-\alpha_{t+1}$. Since this process can be seen as conditional sampling based on the outcome of a biased coin-flip, we refer to the outcome of this flip as the ‘identity’ of the source.
@@ III-A Bounds on the minimax loss
We are now ready to state the main results of this paper. The lower bound on the minimax loss is given in terms of an error term, defined subsequently.
@@
whenever $\sum\limits_{i\in[0:t]}n_{i}\alpha_{i}^{2}\,+h_{t}(\epsilon)\geq k/4$, where $h_{t}(.)$ is defined as in Definition 3.
Definition 3 (Error term for lower bounds).
The error term $h_{t}(\epsilon)$ is defined as
$h_{t}(\epsilon):=\sum_{i=0}^{t}n_{i}\alpha_{i}g_{i}(\epsilon),$
where
$g_{i}(\epsilon)$ $:=\sup_{p}\max_{j}\bar{\mathbf{P}}_{i-1,({p})}\left\{\hat{P}_{i-1}[j]<p[j]-\epsilon\right\}$
denotes the worst case (component-wise and distribution-wise) probability of an $\epsilon$-error in $\hat{P}_{i-1}$ under $\bar{\mathbf{P}}_{i-1,({p})}$ for all $i\neq 0$, and $g_{0}(\epsilon)=0\ \forall\epsilon$.
@@
We can directly observe that when the error term $h_{t}(\nicefrac{{1}}{{4k}})$ is small, the minimax errors are within constant factors of each other; and, as we will see in the baseline, the erasure of the source information leads to a penalty of $\alpha_{t}$ on the effective sample size of batch $t$ (as compared to when the estimators know which distribution each sample is drawn from). As a corollary of the analysis of Theorem 2, we also describe the worst-case error when the sequence of empirical estimators without further processing is used.
@@
This bound most closely resembles the results in [12, 16]; in fact, if each batch is the same size $n$, we find an upper bound of $\nicefrac{{\pi^{2}}}{{6}}\cdot\nicefrac{{1}}{{n}}$, recovering the additional penalty factor of $\nicefrac{{\pi^{2}}}{{6}}$ highlighted therein. We note that this bound increases with $t$, and therefore, in our context, it is perhaps useful only to use the upper bound $\nicefrac{{1}}{{n_{0}}}$ obtained by using the order-optimal estimator for just the $0^{\mathrm{th}}$ batch.
With that said, it is interesting to note that letting the estimate be the empirical distribution of the whole corpus provides a notion of stability ‘for free’—the sum in the RHS of (5) is exactly $\nicefrac{{1}}{{n_{0}}}$ plus a (lower) Riemann sum of the function $\nicefrac{{1}}{{x^{2}}}$ from $n_{0}$ to $\sum_{i=0}^{t}n_{i}$ with intervals $n_{i}$, and therefore is always bounded above by a constant smaller than $1$.
@@ III-B Baseline: oracle-assisted minimax loss
To understand the impact of not knowing the source identity on the minimax loss, it is useful to compare the upper and lower bounds presented above to the loss in the case where the identity of the source of each sample is known in the form of an auxiliary random vector $W:=W_{i}^{n_{i}}\sim(\mathrm{Bern}(\alpha_{i}))^{n_{i}},i\in\{1,\dots,t\}$ provided to the estimator by an oracle such that, conditioned on whether $W_{i}[s]$ is 0 or $1$, the sample $X_{i}[s]$ is distributed either as $p$ or $\hat{P}_{i-1}$ respectively.
@@
Remark 6 (Minimax bounds for other losses).
The bounds in this work have been shown for the $\ell_{1}$ and $\ell_{2}^{2}$ risks, with a focus on the $\ell_{2}^{2}$ risk for clarity of exposition. The $\ell_{1}$ loss is related to other losses used in the model collapse literature via standard inequalities.
  • (Total variation and Wasserstein-1 losses): It is well-known that the total variation distance between two distributions is exactly twice their $\ell_{1}$ distance. In the discrete setting, the Wasserstein $1$-distance is equal to the total variation under the natural metric $\rho(.,.):=\mathds{1}\{.\neq.\}$.
  • (KL divergence and cross-entropy losses): Pinsker’s inequality relates the total variation to the KL divergence. Additionally, the cross-entropy loss/log-loss $E_{p}[-\log q(X)]$ used widely in the generative machine learning literature is exactly equal to $D(p\|q)+H(p)$, where $H(.)$ is the entropy of the true distribution.
  • (Hellinger distance): Le Cam’s inequalities [23, Lemma 2.3]
    $\frac{1}{2}\mathcal{H}^{2}(p,q)\leq\|p-q\|_{TV}\leq\mathcal{H}(p,q)\sqrt{1-\frac{\mathcal{H}^{2}(p,q)}{4}}$
    relate the total variation distance and the Hellinger distance $\mathcal{H}(.,.)$ respectively.
We can therefore directly use the upper and lower bounds provided in this work for the minimax Hellinger distance and the lower bounds for the KL divergence and cross-entropy losses. Using these inequalities, we have lower and upper bounds on the minimax Hellinger distance scaling as the minimax $\ell_{1}$ distance and its square root respectively; and the lower bound on the minimax KL divergence (and therefore also the minimax excess cross-entropy loss) scaling as the square of the minimax $\ell_{1}$ loss.
@@ IV Discussion
It is evident from the bounds shown above that if the mixing factor $\alpha_{t}$ is constant (or even $\Omega(1)$), we can show that the base and oracle-assisted minimax losses are within constant factors of each other; the case where the probability converges to zero is more interesting. This corresponds to cases where the rate of synthetic sampling outpaces the real sampling rate by a considerable margin. Such cases might arise when it is significantly cheaper to sample the synthetic distribution, perhaps also exacerbated by the fact that the estimates get closer to the real distribution with every new batch. The behavior of the bounds in such cases depends on the rates of growth of $\sum_{i=0}^{t}n_{i}\alpha_{i}$ and $\sum_{i=0}^{t}n_{i}\alpha_{i}^{2}$. The absence of source information at the estimator leads to a penalty factor of roughly $\max(\alpha_{i},g_{i-1}(\nicefrac{{1}}{{4k}}))$ in the ‘effective sample size’ of each batch; we show that this can lead to large significant gaps between the base and oracle-assisted losses. Depending on whether or not the $\alpha_{i}$’s are comparable to the $g_{i-1}(\nicefrac{{1}}{{4k}})$’s, there might also be a gap between the guarantees provided by the upper and the lower bounds. In the sequel, we comment on the impact of the error term on the behavior of the bounds and the restrictions on the order-optimal estimator, show cases where differences in the aforementioned rates of growth imply gaps between the base and oracle-assisted losses (suggesting a change in scaling laws), and conclude with a summary of the regimes where it is possible to comment precisely on the minimax error.
@@ IV-A The error term $h_{t}({\nicefrac{{1}}{{4k}}})$
The multi-stage lower bounds depend on the performance of estimators in previous stages through the probabilities $g_{i}\left(\nicefrac{{1}}{{4k}}\right)$. Intuitively, the quantities $g_{i}(\nicefrac{{1}}{{4k}})$ capture the worst-case deviation probability of any component of $\hat{P}_{i-1}$ from $p$ by more than $\nicefrac{{1}}{{4k}}$. Since we focus consider cases where $\hat{P}_{t}$ converges to $p$, this probability is expected to decay to zero. It is perhaps surprising, then, to also observe that the lower bounds decrease as the probabilities $g_{i}(\nicefrac{{1}}{{4k}})$ increase—suggesting that it might be possible to improve the performance of future estimators by estimating badly in the present. The source of this apparent improvement is the loosening of the upper bound on the KL divergence in (16) in instances when the previous estimators perform badly. This suggests that it might be easier for an estimator to perform well when the synthetic samples are from a distribution sufficiently removed from the real one; the sequence of estimators in Theorem 2 does not share this behavior.
We note that the error term is small compared to the other term in the denominator, $\sum_{i=0}^{t}n_{i}\alpha_{i}^{2}$, in quite a lot of cases. Using Chebyshev’s inequality (as demonstrated in the upcoming examples) leads to an upper bound on each of these probabilities in terms of the minimax error $r_{i-1,k}^{\ell}$; the upper and lower bounds thus match whenever $n_{i}\alpha_{i}^{2}$’s are larger than or comparable to $1/\sum_{j=0}^{i-1}n_{j}\alpha_{j}^{2}$. We also note that it is necessary to use Chebyshev’s inequality—as opposed to stronger concentration results such as Hoeffding-style bounds that rely on the specific properties, such as subgaussianity in the case of the Hoeffding bound, of the estimators—to derive a universal upper bound on these probabilities for any sequence of estimators that achieves order-optimal performance.
@@ IV-B The order-optimal estimator
The sequence of estimators in Theorem 2 do not always return distributions; indeed, an assumption on the sample sizes and mixing factors is required to make sure that the outputs stay in $\Delta_{k}$ at every stage. Each estimate in the sequence is also kept unbiased and uncorrelated from the previous estimate; this helps keep the analysis from becoming almost prohibitively cumbersome. A careful analysis of a more general class of nonlinear estimators might help bridge the gap with the lower bounds, proving the error term necessary, or drop the assumptions required. Alternatively, a tighter analysis bounding the aforementioned divergence might lead to a lower bound that matches the upper bound wherever valid.
For the rest of this paper, we analyze the lower bounds assuming ‘good’ estimators—estimators with worst-case loss within at most a constant factor of the lower bound—exist, and are chosen at every stage. In quite a few cases, as it turns out, we find that the error term is small as compared to the rest of the denominator, and the analysis is tight if optimal or near-optimal estimators are assumed to have been used. We now show the main results in action via two examples, one where the bounds match and the error term can be ignored, and one where the error term dominates, but still leads to a gap w.r.t. the oracle-assisted loss.